OpenStudy (anonymous):

Differentiate the function y = 2x*log[10](sqrt(x))

OpenStudy (anonymous):

y' = [(2x)/(ln(10)*sqrt(x)] + [2*log[10](sqrt(x))]

OpenStudy (anonymous):

correct so far?

OpenStudy (anonymous):

or no the derivative of log[10](sqrt(x)) would be (1/[sqrt(x)ln(10)])*(1/[2sqrt(x)])

OpenStudy (anonymous):

no it's not

OpenStudy (anonymous):

well according to the textbook the answer is pretty different

OpenStudy (anonymous):

alright well i've got 1/ln(10) + 2*log[10](sqrt(x))

OpenStudy (anonymous):

that's not final though so help me from there

OpenStudy (anonymous):

somehow the second term (after the +) needs to equal log[10](x) does that make sense

OpenStudy (experimentx):

isn't it something like this http://www.wolframalpha.com/input/?i=d%2Fdx%28+2x*log+base+10+%28sqrt%28x%29%29%29

OpenStudy (anonymous):

sure i guess it's something like that bro jesus you're a great help

OpenStudy (anonymous):

but that's still not the textbook answer

OpenStudy (experimentx):

let me test, dy/dx = 2{dx/dx*log[10](sqrt(x))+x*d(log[10](sqrt(x)))/dx} = 2{log[10](sqrt(x))+x*d(ln(sqrt(x)))/dx*1/ln10} = 2{log[10](sqrt(x))+x*1/sqrt(x)*1/2*1/sqrt(x)*1/ln10} = 2{log[10](sqrt(x))+1/2ln10} my answer.

OpenStudy (anonymous):

alright here look the textbook answer is 1/ln(10) + log[10](x).

OpenStudy (experimentx):

i got almost same answer ... except 1/ln(10) + 2log[10](x)

OpenStudy (anonymous):

i got as far as 1/ln(x) + 2(log[10](x)), the latter of which = 2[log(sqrt(x))/log(10)], so [2*log(sqrt(x))] / log(10) = log(sqrt(x))^2/log(10) ?? which = log(x)/log(10)? which = log[10](x) ???? is that correct? does that make sense? i'm reeeal rusty on the log business

OpenStudy (anonymous):

nah bro according to the product rule that other end of the + sign is gonna be the derivative of 2x * the log[10](sqrt(x)). unless that's not the case. but i'm pretty sure it is. don't see any way around it really. and you got the derivative of 2x multiplying the other factor which is supposed to be left as is, so.... you tell me..... if and how i'm wrong.....

OpenStudy (experimentx):

use this property of log to convert it into natural log. \[ \log_ax = \frac{\ln x}{\ln a} \] I think there's a better example at wikipedia.

OpenStudy (anonymous):

i know that, and it still doesn't explain how 2*log[10](sqrt(x)) = log[10](x).

OpenStudy (anonymous):

oh unless... that equals log[10]([sqrt(x)]^2) ???? is that it?

OpenStudy (anonymous):

oh so it is